Markus Company’s common stock sold for $3.25 per share at the end of this year. The company paid a common stock dividend of $0.65 per share this year. It also provided the following data excerpts from this year’s financial statements: Ending Balance Beginning Balance Cash $ 39,000 $ 35,200 Accounts receivable $ 72,000 $ 56,200 Inventory $ 61,900 $ 72,000 Current assets $ 172,900 $ 163,400 Total assets $ 531,000 $ 550,400 Current liabilities $ 69,000 $ 60,000 Total liabilities $ 156,000 $ 140,400 Common stock, $1 par value $ 132,000 $ 132,000 Total stockholders’ equity $ 375,000 $ 410,000 Total liabilities and stockholders’ equity $ 531,000 $ 550,400 This Year Sales (all on account) $ 850,000 Cost of goods sold $ 493,000 Gross margin $ 357,000 Net operating income $ 173,375 Interest expense $ 9,500 Net income $ 114,712 11. What is the company’s operating cycle?
Answer #1

Operating cycle = Days' Sales of Inventory + Days' Sales Outstanding = 49.57 + 27.53 = 77.10 days

Days’ sales of inventory = 365*average inventory / cost of goods sold = 365*((61900+72000)/2)/493000 = 49.57 days

Days' Sales Outstanding = 365*average accounts receivable / sales = 365*((72000+56200)/2)/850000 = 27.53 days